OpenStudy (anonymous):

I guess i do have a question.. Solve the linear system using substitution 4x-y=5 4x+4y=-4 Been working on it for while..feel like such an idiot.

OpenStudy (amistre64):

do you need to do substitution? cause theres a quicker way

OpenStudy (anonymous):

I know there is...stupid algebra 1 Sucks..

OpenStudy (anonymous):

(I have to)

OpenStudy (amistre64):

define the first equation in terms of y(x)

OpenStudy (amistre64):

4x-y=5 4x-y(x)=5 -4x -4x ------------- -y(x) = 5-4x ; *-1 y(x) = 4x-5

OpenStudy (amistre64):

4x+4y(x)=-4 ; /4 x+y(x)=-1 ; sub in y(x) x+(4x-5) = -1

OpenStudy (anonymous):

Thanks. i can look this over and use it as a reference to other questions. (Gave you the best response..as your the only one who did)

OpenStudy (amistre64):

once you determine a value for "x" use it to define y(x) ... alos known as just "y"

OpenStudy (anonymous):

Thanks a ton.

OpenStudy (amistre64):

good luck ;)